17. One leg of a right triangle is 15. The hypotenuse is 17. What is the perimeter of the triangle?: *
8
40
127.5
578
18. Juan and Anita are standing together in the schoolyard. Juan, who is 6 feet 2 inches tall, casts a shadow that is 54 inches long. At the same time, Anita casts a shadow that is 46 inches long. How tall is Anita? (Round your answer as necessary): *
86.4 in.
5 ft 3 in.
5 ft 1 in.
64 in.
Please respond to the question above.
19. Factor: 5x2 – 13xy – 6y2: *
x(5x-13y) - y(13x+6y)
(5x+2y)(x-3y)
(5x-3y)(x+2y)
The expression cannot be factored.
Please respond to the question above.
20.: *

Choice a
Choice b
Choice c
Choice d
Please respond to the question above.
21. A boat cruises downstream for 2 hours before heading back. It takes 3.5 hours going upstream to get back. If the speed of the stream is 8 mph, what is the speed of the boat in still water?: *
14.7 mph
19.6 mph
24.8 mph
29.3 mph
Please respond to the question above.
22. Complete the following synthetic division: *

3 -15 6 -27
1 -8 26 -87
1 -2 -4 -21
4 7 23 60
Please respond to the question above.
23. Which of the following is a correct classification of the number: *

Irrational
Rational
Integer
Natural
Please respond to the question above.
24. At what point do the lines y = 4x + 7 and y = -4x – 11 intersect?: *
(-5/4, 2)
(9/4, 16)
(-9/4, -2)
The lines do not intersect.
Please respond to the question above.
25. Solve |3x+1| < -5: *
x < -2
-2 < x < 4/3
all real numbers
no solution
Please respond to the question above.

Answer by edjones(8007)

17)
15^2+b^2=289
b^2=64
b=8
8+15+17=40 perimeter.
.
Ed